Test the following series for convergence 

∑
n=1
∞ [✓n^4+9 -✓n^4-9]
1
Expert's answer
2021-06-16T17:22:04-0400

"\sqrt{n^4 +9}-\sqrt{n^4 -9}\\
\text{Let} 
a _n=\sqrt{n^4 +9}-\sqrt{n^4 -9}\\
\text{Multiplying and dividing }
a _n
\text{by }
\sqrt{n^4 +9}+\sqrt{n^4 -9}.\\
a_n=\frac{18}{\sqrt{n^4 +9}+\sqrt{n^4 -9}}\\
Then,\\
\frac{18}{\sqrt{n^4 +9}+\sqrt{n^4 -9}}=
\frac{18}{n^2(\sqrt{1 +\frac{9}{n^4}}+\sqrt{1-\frac{9}{n^4}})}\leq\frac{18}{n^2}
\\
\text{Let} \\
b_n=\frac{18}{n^2}.\\
\text{Since, for all n,}
0≤a_n≤b_n.\\
and ∑ _{i=1}^∞b_n =∑ _{i=1}^∞\frac{18}{n^2}
(\text{Converges by p-test})\\
\text{By comparision test,}\\
\sum a_n \text{converges.}"